OpenStudy (anonymous):

sec(pi + theta) is negative right? @ganeshie8

ganeshie8 (ganeshie8):

we got a doubt, so lets derive it and see... :)

ganeshie8 (ganeshie8):

\[\sec (\pi + \theta) = \dfrac{1}{\cos(\pi + \theta)}\]

OpenStudy (anonymous):

okay! :)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

use angle sum identity for the denominator

ganeshie8 (ganeshie8):

\[\sec (\pi + \theta) = \dfrac{1}{\cos(\pi + \theta)} = \dfrac{1}{\cos \pi \cos \theta - \sin \pi \sin \theta}\]

OpenStudy (anonymous):

yess

ganeshie8 (ganeshie8):

whats the value of \(\sin \pi\) ?

ganeshie8 (ganeshie8):

whats the value of \(\cos \pi\) ?

OpenStudy (sheraz12345):

Negative as it lies in third quadrant

OpenStudy (anonymous):

sin 180 = 0 cos 180 = -1

OpenStudy (anonymous):

yes i think it would be negative as it lies in third quadrant where sec is negative

ganeshie8 (ganeshie8):

\[\sec (\pi + \theta) = \dfrac{1}{\cos(\pi + \theta)} = \dfrac{1}{\cos \pi \cos \theta - \sin \pi \sin \theta} = \dfrac{1}{-1 \cos \theta - 0\sin \theta} \]

OpenStudy (sheraz12345):

Yes @No.name

ganeshie8 (ganeshie8):

\[= \dfrac{1}{ \cos \theta } = -\sec \theta \]

ganeshie8 (ganeshie8):

thats a good observation ! if all we want to know is the polarity

OpenStudy (anonymous):

thanks , well i found a way in my book to remember it , i will post it for others to see

ganeshie8 (ganeshie8):

these are always tricky, would love to see it :)